# Hiding Functions within Functions: Steganography by Implicit Neural Representations

### v1

@@ III-A Representing A Single Image With A Function
where the sum is over all pixel positions. It is worth noting that the representation $f_{\theta}$ does not depend on the number of pixels. Therefore, unlike most image representations, the representation $f_{\theta}$ is resolution-independent [22] [23] .
III-B Representing General Data With Functions
This example for images can be generalized to more general data. Let $x\in X$ represent the coordinates and $y\in Y$ represent the features, and suppose we have a set of data points with pairs of coordinates and features $\{(x_{i},y_{i})\}_{i=1}^{n}$. For example, for an image, $x=(x;y)$ corresponds to the pixel positions, and $y=(r;g;b)$ corresponds to the RGB values, and $\{(x_{i},y_{i})\}_{i=1}^{n}$ corresponds to the set of all pixel positions and RGB values. Given a set of coordinates and their corresponding features, we can learn a function $f_{\theta}:X\rightarrow Y$ to represent this data by minimizing the following loss function:
@@ III-B Representing General Data With Functions
Figure 2: Steganography Framework based on Implicit Neural Representations.
Recent research has shown that learning function representations by minimizing (1) results in a bias towards low-frequency functions [8] [23]. While several methods have been proposed to mitigate this issue, we use the Random Fourier Features (RFF) encoding introduced by Tancik et al. [24] as it does not bias towards axial changes [8] and does not require specialized initialization [23]. Specifically, for a given coordinate $x\in{{R}^{d}}$, the encoding function $\gamma:{{R}^{d}}\rightarrow{{R}^{2m}}$is defined as follows:
@@
where $B$ is a random matrix of size $m\times d$ (possibly learnable) whose elements are typically sampled from $N(0,\sigma^{2})$. The number of frequencies $m$ and the variance of the elements of matrix $B$, $\sigma^{2}$, are hyperparameters. To learn high-frequency functions, we simply encode $x$ before passing it through the MLP $f_{\theta}$, i.e., $f_{\theta}(\gamma(x))$, and then minimize (1). Using an MLP with ReLU activation to learn the function representation of an image fails to capture high-frequency details, while using RFF encoding followed by a ReLU MLP allows accurate reproduction of the image.
IV Problem Formulation
IV-A Function Hiding
For simplicity, let’s use the implicit representation of images as an example to describe our method, as shown in Fig 2. Assuming the secret message to be transmitted is denoted as $m$, for the sender, the following three steps need to be completed sequentially to achieve steganography:
@@ IV-A Function Hiding
In the first step, as described in (4), we parameterize the image $m$, where ${{f}_{\theta}}$ represents the initial function before parameterization, ${{f}_{\theta}^{m}}$ represents the parameterized secret function, and $IN{{R}_{p}}$ denotes the process of implicit representation.In the second step, as in (5), the sender hides ${{f}_{\theta}^{m}}$ into another function ${{f}_{\phi}}$. We introduce a key $k$ to build the stego function ${{f}_{\phi}}$ based on ${{f}_{\theta}^{m}}$. In this paper, the key $k$ is used to control the function structure of ${{f}_{\phi}}$ and the location of ${{f}_{\theta}^{m}}$.In the third step, as in (6), we utilize ${{f}_{\phi}}$ to achieve the implicit representation of a specific cover image $c$, obtaining the final stego function ${{f}_{\phi}^{c}}$, where $\phi$ is the parameter set of the stego function ${{f}_{\phi}^{c}}$.
For the message receiver, it is only necessary to use the secret key $k$ and the secret function ${{f}}_{\phi}^{c}$ obtained from the public channel to recover the secret message $m$. The receiver first recovers the secret function ${{f}}_{\theta}^{m}$ using (7), and then utilizes the sampling process of INR to recover the secret message $m$ as follows:
@@
IV-B Data Representation
From the principles of implicit representation, it can be seen that in order to train the parameters of this function, the data must be transformed into a set of coordinates and features, i.e., $m=\{({{x}}_{i},{{y}}_{i})\}_{i=1}^{n}$. Such a set of coordinates and features corresponds to the input/output pairs of the function, allowing us to learn the function’s parameters. An individual sample point corresponds to a set of coordinates and features (e.g., an image is a set of position coordinates and individual pixel values). The use of a set representation of coordinates and features is very flexible, as it is independent of whether the data comes from a grid or the sampling resolution. The key is that constructing the problem entirely based on sets allows us to divide individual data points into subsets and train on them. Specifically, given an individual data point $s=\{({{x}}_{i},{{y}}_{i})\}_{i=1}^{n}$, such as a set of n pixels, we can randomly sample k elements, for example, select $m$ pixels from the $n$ pixels of the entire image. Training on these subsets eliminates the direct dependency on data resolution. By adopting this approach, it is easy to obtain an implicit expression network that represents low-resolution data when higher resolution data is not required. Similarly, when training 3D shapes, there is no need to pass the entire voxel grid to the model. Instead, training can be done on subsets of the voxel grid, saving a significant amount of memory. This is impossible in standard convolutional models, as they are directly associated with the grid resolution. Furthermore, the training method based on the set representation of coordinates and features allows us to model more peculiar data, such as function distributions on manifolds. In fact, as long as we can define a coordinate system on the manifold (e.g., polar coordinates on a sphere), our steganographic method can be applied to all forms of data, whether it is used as secret messages or cover data.
V The Proposed Method
V-A Secret Function
In order to keep our method as simple as possible, we adopt MLP as the underlying network for implicit representation. The weight between the $l$-th and $(l+1)$-th layers can be written in matrix form as $W_{mn}^{l}$, and the bias can be written in vector form as the bias vector $b$. The dimensions of the weight matrix are $m\times n$, where $m$ represents the number of neurons on the right side of the two layers, and $n$ represents the number of neurons on the left side. The $i$-th row and $j$-th column of the matrix represent the weight between the $i$-th neuron in the right layer and the $j$-th neuron in the left layer. Thus, the equation can be simplified as:
@@ V-A Secret Function
For a given secret message $m$, we use (1) to learn the parameters of the ${{f}_{\theta}}$. The resulting function representation is denoted as ${{f}_{\theta}^{m}}=\{{{P}_{\theta}},{{S}_{\theta}}\}$, referred to as the secret function.
Figure 3: Three Strategies for Constructing Stego Functions
V-B Stego Function
@@
More importantly, we want the expanded stego function to have certain functionality, making it more concealable during transmission, i.e., using some behavior to conceal the steganographic behavior. Typically, we have two choices. One is to use this stego function (network) to perform a regular machine learning tasks, such as image classification, as done in [49] [50]. In this paper, we adopt another method that is consistent with the secret message, namely using the implicit representation to conceal the steganographic behavior. When both the secret function and the stego function are used to represent the same type of data, taking images as an example, we design the following three strategies for function expansion, as shown in Fig 3.
(1) Horizontal Expansion. This strategy refers to inserting new layers after the secret function to construct a new function, while keeping the previous network parameters and structure unchanged, as shown in Fig 3.(a). (2) Vertical Expansion. Another method of expanding the function is to keep the number of layers unchanged. Since both the cover function and the secret function are used to represent images, we only need to increase the number of hidden units in the hidden layers, except for the input layer and the final output layer, as shown in 3(b). (3) Mixed Expansion. Mixed expansion involves both vertical and horizontal expansions, i.e., increasing the number of layers and the number of neurons in existing hidden layers at the same time. This is a general operation, as shown in 3(c), where all the content, parameters, and structure of the secret network are hidden within a single network.
@@
When the secret function and the cover function are used to represent different types of data, we can naturally adopt the mixed expansion for stego function construction. For example, hiding the implicit representation of an image within the implicit representation of a 3D model. Once the structure of the stego function $f_{\phi}$ is determined, we can train it to represent a specific cover data $c$. It should be noted that the common characteristic of these three expansion methods is that the secret function $f^{m}_{\theta}$, remains intact within the stego function, and the dependency between layers is not disrupted. This will assist us in designing the recovery of the secret function.
Since the structure of the cover function needs to be determined before expansion, we can consider the function expansion operation as relying on the shared secret information $k$, as shown in (5). In our scheme, the key $k=\{k_{1},k_{2},...,k_{L}\}$ indicates the positions of $f^{m}_{\theta}$ within $f_{\phi}$, where each layer corresponds to a binary stream $k_{l}$ of $d_{k_{l}}$ length, representing the number of neurons in that layer. Each bit corresponds to a neuron, where 0 represents the position of the masked neuron, and 1 represents the position of the secret neuron. For example, the key $k$ corresponding to the stego function $f_{\phi}$ in 3(c) is $\{{00,1010,1010101,1101,000\}}$.
@@
V-C Training of the Stego Function
We use (1) to achieve the implicit representation of the cover image $c$ by the stego function $f_{\phi}^{c}$. $\phi$ includes two different types of parameters: the original parameters $\theta$ of the secret function $f_{\theta}^{m}$, and the parameters $\phi_{e}$ of the expanded neurons. To ensure that we can extract $f_{\theta}^{m}$ from $f_{\phi}^{c}$ without loss, during the training process, we fix the parameters $\theta$ of $f_{\theta}^{m}$ and only adjust the parameters $\phi_{e}$ of the expanded neurons. To do this, we introduce a mask $M^{k}$ to partially optimize $f_{\phi}$. $M^{k}$ is a binary mask vector $v^{k}$ with the same size as the parameter set $\phi$ of $f_{\phi}$,
@@ V-C Training of the Stego Function
Let $\eta$ be the learning rate and $\odot$ denotes element-wise multiplication. We update $f_{\phi}^{c}$ using gradient descent as follows:
@@
V-D Secret Function Recovery
@@
Figure 4: Secret Function Recovery.
As shown in Fig 4, the receiver first overlays the values of $k$ onto the neuron nodes $n$ of the stego function $f_{\phi}^{c}$, i.e., if $k^{l}_{i_{l}}=1$, then $n(i_{l})^{l}=1$. , where $n(i_{l})^{l}$ represents the $i_{l}$-th node on the $l$-th layer. We only need to keep the weights ${{w}_{{{i}_{l-1}},{{i}_{l}}}^{l-1,l}}$ that satisfied both $n(i_{l})^{l}=1$ and $n(i_{l-1})^{l-1}=1$. Then we can obtain the weight matrix of layer $l$ in $f_{\theta}^{m}$ as $W^{\theta}_{l,m\times n}$, where $m$ and $n$ represent the dimensions of the weight matrix. Thus, the parameters of the secret function $f_{\theta}^{m}$ can be represented as:
@@
Finally, by counting the number of 1s in each layer of $k$, which represents the number of secret neurons, we can obtain the model structure:
@@
V-E Secret Message Extraction
To extract the secret message $m$ from the recovered secret function $f^{\theta}_{m}$, it is essentially a problem of sampling from the implicit representation function, i.e.,
@@ V-E Secret Message Extraction
VI Experiment
VI-A Setup
We evaluate our model on various resolution images from CelebAHQ [25] Div2k dataset [54] and COCO dataset [55] , and climate data from the ERA5 dataset [27] . We use an MLP for secret function and stego function to represent the secret message and the cover data. Our model is implemented using PyTorch [28] , and we utilize the implicit representation framework provided in [29] . All training is conducted on a single 2080Ti GPU with 11GB memory.
@@ VI-A Setup
Figure 5: A group of experimental results on the CelebA-HQ database.
@@
VI-B Performance
VI-B1 Representation of The Secret Message
For the secret image $m$, we use the Peak Signal-to-Noise Ratio (PSNR) of the original image and the image obtained by sampling the secret function to measure the implicit representation capability. Fig 6 shows the representation results of images with different resolutions in the case of an MLP network structure of [64, 64, 64]. From left to right, the resolutions are 64x64, 128x128, 256x256, 512x512, and 1024x1024, repectively, in which the 64x64 image comes from CelebAHQ dataset, the 128x128 and 256x256 images come from COCO dataset, and the 512x512 and 1024x1024 images come from the Div2K dataset.
Figure 6: PSNR of images sampled by an implicit representation network at different resolutions.
Figure 7: PSNR Curve of Images Obtained at Different Resolutions.
@@ VI-B1 Representation of The Secret Message
Fig 7 shows the PSNR values of the secret image and the sampled secret image from the secret function at different resolutions. It can be seen from the figure that as the resolution increases, the PSNR decreases to some extent. This is because the method of calculating differences only treats the image as isolated pixels and ignores some of the visual features contained in the image contents.
VI-B2 Recoverability
TABLE I: The BER for Different Structures of Secret Functions

### v2

@@ 3.1 Representing an Image with Function
where the sum is over all pixel positions. It is worth noting that the representation $f_{\theta}$ does not depend on the number of pixels. Therefore, unlike most image representations, the representation $f_{\theta}$ is resolution-independent [46, 47].
3.2 Representing General Data With Functions
This example for images can be generalized to more general data. Let $x\in X$ represent the coordinates and $y\in Y$ represent the features, and suppose we have a set of data points with pairs of coordinates and features $\{(x_{i},y_{i})\}_{i=1}^{n}$. For example, for an image, $x=(x;y)$ corresponds to the pixel positions, and $y=(r;g;b)$ corresponds to the RGB values, and $\{(x_{i},y_{i})\}_{i=1}^{n}$ corresponds to the set of all pixel positions and RGB values. Given a set of coordinates and their corresponding features, we can learn a function $f_{\theta}:X\rightarrow Y$ to represent this data by minimizing the following loss function as described in Eq. (2).
@@ 3.2 Representing General Data With Functions
Fig.  2: Steganography Framework based on Implicit Neural Representations.
Recent research has shown that learning function representations by minimizing Eq. (1) results in a bias towards low-frequency functions[9, 47]. While several methods have been proposed to mitigate this issue, we use the Random Fourier Features (RFF) encoding introduced by Tancik [48] as it does not bias towards axial changes [9] and does not require specialized initialization [47]. Specifically, for a given coordinate $x\in{{R}^{d}}$, the encoding function $\gamma:{{R}^{d}}\rightarrow{{R}^{2m}}$ is defined as follows:
@@
where $B$ is a random matrix of size $m\times d$ (possibly learnable) whose elements are typically sampled from $N(0,\sigma^{2})$. The number of frequencies $m$ and the variance of the elements of matrix $B$, $\sigma^{2}$, are hyperparameters. To learn high-frequency functions, we simply encode $x$ before passing it through the MLP $f_{\theta}$, i.e., $f_{\theta}(\gamma(x))$, and then minimize Eq. (1). Using an MLP with ReLU activation to learn the function representation of an image fails to capture high-frequency details, while using RFF encoding followed by a ReLU MLP allows accurate reproduction of the image.
4 Problem Formulation
4.1 Function Hiding
For simplicity, let’s use the implicit representation of images as an example to describe our method, as shown in Fig. 2. Assuming the secret message to be transmitted is denoted as $m$, for the sender, the following three steps need to be completed sequentially to achieve steganography:
@@ 4.1 Function Hiding
In the first step, as described in Eq. (4), we parameterize the image $m$, where ${{f}_{\theta}}$ represents the initial function before parameterization, ${{f}_{\theta}^{m}}$ represents the parameterized secret function, and $IN{{R}_{p}}$ denotes the process of implicit representation. In the second step, as in Eq. (5), the sender hides ${{f}_{\theta}^{m}}$ into another function ${{f}_{\phi}}$. We introduce a key $k$ to build the stego function ${{f}_{\phi}}$ based on ${{f}_{\theta}^{m}}$. In this paper, the key $k$ is used to control the function structure of ${{f}_{\phi}}$ and the location of ${{f}_{\theta}^{m}}$. In the third step, as in Eq. (6), we utilize ${{f}_{\phi}}$ to achieve the implicit representation of a specific cover image $c$, obtaining the final stego function ${{f}_{\phi}^{c}}$, where $\phi$ is the parameter set of the stego function ${{f}_{\phi}^{c}}$.
For the message receiver, it is only necessary to use the secret key $k$ and the secret function ${{f}}_{\phi}^{c}$ obtained from the public channel to recover the secret message $m$. The receiver first recovers the secret function ${{f}}_{\theta}^{m}$ using Eq. (7), and then utilizes the sampling process of INR to recover the secret message $m$ as follows:
@@
4.2 Data Representation
From the principles of implicit representation, it can be seen that in order to train the parameters of this function, the data must be transformed into a set of coordinates and features, i.e., $m=\{({{x}}_{i},{{y}}_{i})\}_{i=1}^{n}$. Such a set of coordinates and features corresponds to the input/output pairs of the function, allowing us to learn the function’s parameters. An individual sample point corresponds to a set of coordinates and features (i,e., an image is a set of position coordinates and individual pixel values). The use of a set representation of coordinates and features is very flexible, as it is independent of whether the data comes from a grid or the sampling resolution. The key is that constructing the problem entirely based on sets allows us to divide individual data points into subsets and train on them. Specifically, given an individual data point $s=\{({{x}}_{i},{{y}}_{i})\}_{i=1}^{n}$, such as a set of n pixels, we can randomly sample k elements(i.e., select $m$ pixels from the $n$ pixels of the entire image).
Training on these subsets eliminates the direct dependency on data resolution. By adopting this approach, it is easy to obtain an implicit representation network that represents low-resolution data when higher resolution data is not required. Similarly, when training 3D shapes, there is no need to pass the entire voxel grid to the model. Instead, training can be done on subsets of the voxel grid, saving a significant amount of memory. This is impossible in standard convolutional models, as they are directly associated with the grid resolution. Furthermore, the training method based on the set representation of coordinates and features allows us to model more peculiar data, such as function distributions on manifolds. In fact, as long as we can define a coordinate system on the manifold (i.e., polar coordinates on a sphere), our steganographic method can be applied to all forms of data, whether it is used as secret messages or cover data.
5 The Proposed Method
5.1 Secret Function
In order to keep our method as simple as possible, we adopt MLP as the underlying network for implicit representation. The weight between the $l$-th and $(l+1)$-th layers can be written in matrix form as $W_{mn}^{l}$, and the bias can be written in vector form as the bias vector $b$. The dimensions of the weight matrix are $m\times n$, where $m$ represents the number of neurons on the right side of the two layers, and $n$ represents the number of neurons on the left side. The $i$-th row and $j$-th column of the matrix represent the weight between the $i$-th neuron in the right layer and the $j$-th neuron in the left layer. Thus, the equation can be simplified as described in Eq. (9) and Eq. (10).
@@ 5.1 Secret Function
For a given secret message $m$, we use Eq. (1) to learn the parameters of the ${{f}_{\theta}}$. The resulting function representation is denoted as ${{f}_{\theta}^{m}}=\{{{P}_{\theta}},{{S}_{\theta}}\}$, referred to as the secret function.
Fig.  3: Three Strategies for Constructing Stego Functions
5.2 Stego Function
@@
More importantly, we want the expanded stego function to have certain functionality, making it more concealable during transmission (i.e., using some behavior to conceal the steganographic behavior). Typically, we have two choices. One is to use this stego function (network) to perform a regular machine learning tasks, such as image classification, as done in [7, 8]. In this paper, we adopt another method that is consistent with the secret message, namely using the implicit representation to conceal the steganographic behavior. When both the secret function and the stego function are used to represent the same type of data, taking images as an example, we design the following three strategies for function expansion, as shown in Fig. 3.
(1) Horizontal Expansion. This strategy refers to inserting new layers after the secret function to construct a new function, while keeping the previous network parameters and structure unchanged, as shown in Fig. 3 (a).
(2) Vertical Expansion. Another method of expanding the function is to keep the number of layers unchanged. Since both the cover function and the secret function are used to represent images, we only need to increase the number of hidden units in the hidden layers, except for the input layer and the final output layer, as shown in Fig. 3 (b).
(3) Mixed Expansion. Mixed expansion involves both vertical and horizontal expansions, i.e., increasing the number of layers and the number of neurons in existing hidden layers at the same time. This is a general operation, as shown in Fig. 3 (c), where all the content, parameters, and structure of the secret network are hidden within a single network.
@@
When the secret function and the cover function are used to represent different types of data, we can naturally adopt the mixed expansion for stego function construction. For example, hiding the implicit representation of an image within the implicit representation of a 3D model. Once the structure of the stego function $f_{\phi}$ is determined, we can train it to represent a specific cover data $c$. It should be noted that the common characteristic of these three expansion methods is that the secret function $f^{m}_{\theta}$ remains intact within the stego function, and the dependency between layers is not disrupted. This will assist us in designing the recovery of the secret function.
Since the structure of the cover function needs to be determined before expansion, we can consider the function expansion operation as relying on the shared secret information $k$, as shown in Eq. (5). In our scheme, the key $k=\{k_{1},k_{2},...,k_{L}\}$ indicates the positions of $f^{m}_{\theta}$ within $f_{\phi}$, where each layer corresponds to a binary stream $k_{l}$ of $d_{k_{l}}$ length, representing the number of neurons in that layer. Each bit corresponds to a neuron, where 0 represents the position of the masked neuron, and 1 represents the position of the secret neuron. For example, the key $k$ corresponding to the stego function $f_{\phi}$ in Fig. 3 (c) is $\{{00,1010,1010101,1101,000\}}$.
@@
5.3 Training of the Stego Function
We use (1) to achieve the implicit representation of the cover image $c$ by the stego function $f_{\phi}^{c}$. $\phi$ includes two different types of parameters: the original parameters $\theta$ of the secret function $f_{\theta}^{m}$, and the parameters $\phi_{e}$ of the expanded neurons. To ensure that we can extract $f_{\theta}^{m}$ from $f_{\phi}^{c}$ without loss, during the training process, we fix the parameters $\theta$ of $f_{\theta}^{m}$ and only adjust the parameters $\phi_{e}$ of the expanded neurons. To do this, we introduce a mask $M^{k}$ to partially optimize $f_{\phi}$. $M^{k}$ is a binary mask vector $v^{k}$ with the same size as the parameter set $\phi$ of $f_{\phi}$, as computed in Eq. (11)
@@ 5.3 Training of the Stego Function
Let $\eta$ be the learning rate and $\odot$ denotes element-wise multiplication. We update $f_{\phi}^{c}$ using gradient descent as described in Eq. (12):
@@
5.4 Secret Function Recovery
@@
Fig.  4: Secret Function Recovery.
As shown in Fig. 4, the receiver first overlays the values of $k$ onto the neuron nodes $n$ of the stego function $f_{\phi}^{c}$, (i.e., if $k^{l}_{i_{l}}=1$, then $n(i_{l})^{l}=1$, where $n(i_{l})^{l}$ represents the $i_{l}$-th node on the $l$-th layer). We only need to keep the weights ${{w}_{{{i}_{l-1}},{{i}_{l}}}^{l-1,l}}$ that satisfy both $n(i_{l})^{l}=1$ and $n(i_{l-1})^{l-1}=1$, and then we can obtain the weight matrix of layer $l$ in $f_{\theta}^{m}$ as $W^{\theta}_{l,m\times n}$, where $m$ and $n$ represent the dimensions of the weight matrix. Thus, the parameters of the secret function $f_{\theta}^{m}$ can be represented as described in Eq. (13).
@@
Finally, by counting the number of 1s in each layer of $k$, which represents the number of secret neurons, we can obtain the model structure as described in Eq. (14)
@@
5.5 Secret Message Extraction
To extract the secret message $m$ from the recovered secret function $f^{\theta}_{m}$, it is essentially a problem of sampling from the implicit representation function, as described in Eq. (15).
@@ 5.5 Secret Message Extraction
6 Experiment
6.1 Setup
We evaluate our model on various resolution images from CelebAHQ[49], Div2k [50] and COCO dataset [51], and climate data from the ERA5 dataset[52]. We also use common datasets NeRF Synthetic[9] with forward 360°scenes lego, drums, chair from NeRF-Synthetic. We use an MLP for secret function and stego function to represent the secret message and the cover data. Our model is implemented using PyTorch [53], and we utilize the implicit representation framework provided in [45]. All training is conducted on a single 2080Ti GPU with 11GB memory.
@@ 6.1 Setup
Fig.  5: A group of experimental results on the CelebA-HQ database.
@@
6.2 Performance
6.2.1 Representation of The Secret Message
For the secret image $m$, we use the Peak Signal-to-Noise Ratio (PSNR) of the original image and the image obtained by sampling the secret function to measure the implicit representation capability. Fig. 6 shows the representation results of images with different resolutions in the case of an MLP network structure of [64, 64, 64]. From left to right, the resolutions are 64x64, 128x128, 256x256, 512x512, and 1024x1024, respectively, in which the 64x64 image comes from CelebAHQ dataset, the 128x128 and 256x256 images come from COCO dataset, and the 512x512 and 1024x1024 images come from the Div2K dataset.
Fig.  6: The images sampled by an implicit representation network at different resolutions.
Fig.  7: The PSNR curve of images obtained at different resolutions.
@@ 6.2.1 Representation of The Secret Message
Figure 7 shows the PSNR values of the secret image and the sampled secret image from the secret function at different resolutions. It can be seen from the figure that as the resolution increases, the PSNR decreases to some extent. This is because the method of calculating differences only treats the image as isolated pixels and ignores some of the visual features contained in the image contents.
6.2.2 Recoverability
Table 1: The BER for different structures of secret functions